### v1

@@ 1 Introduction
For meshfree methods, techniques that rely on knowing the exact integral of a class of functions (such as RBFs) on the domain are typically not available, since these exact integrals are often unknown. Meshfree integration on surface point clouds is therefore often limited to Monte Carlo methods, which have been employed for eigenfunction estimation using the weak form of differential operators [10, 35]. Even though the underlying meshfree interpolation technique used is high-order, Monte Carlo integration is not, which limits the overall rate of convergence. Monte Carlo integration also requires knowledge of the distribution from which a point cloud is sampled, which may not be available, and it may not be straightforward to sample points from a surface uniformly. If the points satisfy a minimum distance condition, so that no two points are too close together, a potential integration technique for non-uniform point clouds is to use local projections to the tangent plane and a triangulation to compute Voronoi cell areas (see [16]), though this has a limited order of convergence. Existing high-order meshfree integration methods on surfaces are often restricted to specific surfaces where exact integrals of RBFs are known, such as the sphere [13]. Reeger, Fornberg, and Watts [22] have also introduced a high-order ($\mathcal{O}\left(h^{7}\right)$, where $h$ is the point spacing), RBF-based method for general closed surfaces that requires first obtaining a triangulation of the point cloud. This is essentially a hybrid approach between a meshfree and meshed method in that while a triangulation of the point cloud is required, a meshfree method is ultimately used to compute integrals on each surface patch projected to a triangle with high accuracy. A version for surfaces with boundary has also been developed [23].
@@
Our contribution consists of two super-algebraically convergent methods based on the divergence theorem and norm-minimizing Hermite-Birkhoff interpolation schemes. Such interpolation schemes include Hermite RBFs [26, 7] and minimum norm methods using Fourier series [4, 31]. We review these interpolation methods in Section 2, including a generalized convergence statement in Proposition 2.1, a discussion on a general approach to high-order convergence in 2.2, and two examples in 2.3. We then introduce our integration methods in Section 3. Our approaches do not require specific quadrature points (the points may be randomly scattered and non-uniformly selected) and only require the knowledge of basic domain information such as normal vectors at scattered points on the surface. One method can be used to directly evaluate the ratio of two integrals on a closed surface or a surface with boundary; this is useful when a constant multiple of the integral is all that is needed. For example, the method could be used when the average value of a function on the surface is desired. The second method requires surfaces to have boundary, but we discuss simple strategies for dividing closed surfaces into multiple surfaces with boundary in Subsection 3.5. Unlike the first method, the second method estimates integrals themselves rather than a ratio, which can be used to estimate surface area with high precision. We show that these schemes converge at the same rate as the underlying interpolation methods, regardless of the point distribution; this leads to arbitrarily high-order schemes. Then, we test the integration techniques for a variety of problems on flat domains and on surfaces in Section 4. We conclude by introducing a more general view of norm-minimizing meshfree methods suitable for handling singularities, and then adapt our methods for singular integrals in Subsection 4.2; such integrals appear in boundary integral solutions to PDEs. Code for running the numerical tests in this paper is available at https://github.com/venndaniel/meshfree_integration.
@@ 2.1 Norm-Minimizing Hermite-Birkhoff Interpolants
(3) $\mathcal{G}u\biggr|_{\partial S}=g,$
@@
where $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}\subset\partial S$ is a set of points on the boundary of $S$. The goal is then to ensure that $\mathcal{F}\tilde{u}\to f$ on $S$ and $\mathcal{G}\tilde{u}\to g$ on $\partial S$ in some sense as $\tilde{N},\tilde{N}_{\partial}\to\infty$. If the PDE given by (1) and (3) is suitably well-posed, we may then be able to show that $\tilde{u}\to u$, with the manner of convergence ($L^{2},L^{\infty},H^{p}$, pointwise, etc.) depending on the PDE and method chosen (see, for example, Proposition 11 of [31]).
Of course, simply imposing $\left(\mathcal{F}\tilde{u}\right)\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)$ on each $\boldsymbol{x}_{j}$ is not enough to ensure that $\mathcal{F}\tilde{u}\to f$ as more collocation points are added. A classic counterexample is the interpolation of Runge’s function ($f\left(x\right)=1/\left(1+25x^{2}\right)$) with polynomials of increasing degree on $\left[-1,1\right]$ and evenly spaced points $\left\{\boldsymbol{x}_{j}\right\}$; this corresponds to an identity operator $\mathcal{F}$ and results in an interpolant $\tilde{u}$ that is unbounded as $\tilde{N}\to\infty$. While basic, this counterexample demonstrates one possible manner in which interpolation can fail: the interpolant can become unbounded. This motivates a constrained approach. For some Hilbert space of functions $\mathcal{H}$, we solve
@@
For this problem to be uniquely solvable, we require $\mathcal{L}$ to be bounded in the $\mathcal{H}\to\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ sense, which means that both $\mathcal{F}$ and $\mathcal{G}$ evaluated at individual points must be bounded as operators from $\mathcal{H}$ to $\mathbb{C}$. We also need the constraint set to be non-empty, which is often the case due to the finite number of collocation points even if the PDE given by (1) and (3) is not solvable globally. In this case, problem (6) has a few important properties. Let $\tilde{u}\in\mathcal{H}$ be the solution to (6), then $\tilde{u}\in\mathcal{R}\left(\mathcal{L}^{*}\right)=\left(\mathcal{N}\left(\mathcal{L}\right)\right)^{\perp}$ (noting that $\mathcal{R}\left(\mathcal{L}^{*}\right)$ is finite-dimensional and therefore closed). We can therefore find $\tilde{u}$ by solving
@@
Note that $u_{n}$ is well-defined due to a standard result in functional analysis (see Theorem 3.3-1 of Kreyszig’s text [14], for example), since $V_{n}$ is non-empty, closed, and convex by assumption. Furthermore, $\left\{\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\right\}_{n=1}^{\infty}$ is non-decreasing, since $V_{n_{1}}\supseteq V_{n_{2}}$ whenever $n_{1}<n_{2}$. For any convex combination $\left(1-\alpha\right)\tilde{u}_{n_{1}}+\alpha\tilde{u}_{n_{2}}$ such that $n_{1}<n_{2}$ and $\alpha\in\left(0,1\right)$,
$\left\|u_{n_{1}}\right\|_{\mathcal{H}}^{2}$ $\leq\left\|\alpha\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right)+\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}$
$=\left\|u_{n_{1}}\right\|_{\mathcal{H}}^{2}+\alpha^{2}\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}+2\alpha\,\text{Re}\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}},\tilde{u}_{n_{1}}\right)_{\mathcal{H}}$
@@
Finally, $\left\{\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\right\}_{n=1}^{\infty}$ is bounded and non-decreasing, and therefore converges, which implies $\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}\to 0$ as $n_{1},n_{2}\to\infty$ due to (11). The sequence $\left\{\tilde{u}_{n}\right\}_{n=1}^{\infty}$ is therefore Cauchy and must converge to some $\tilde{u}_{\infty}\in\mathcal{H}$, since $\mathcal{H}$ is a Hilbert space. Furthermore, for each $m$, $\left\{\tilde{u}_{n}\right\}_{n=m}^{\infty}\subset V_{m}$, which is closed, so $\tilde{u}_{\infty}\in V_{m}$ as well. Finally, $\tilde{u}_{\infty}\in\bigcap_{n=1}^{\infty}V_{n}$.
@@
For our optimization problem applied to PDEs, assume that $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\infty}$ is dense in $S$ and that $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\infty}$ is dense in $\partial S$. Then, Corollary 2.3 implies that if our solution $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}$ to the optimization problem (6) remains bounded in $\mathcal{H}$ as $\tilde{N},\tilde{N}_{\partial}\to\infty$, there is at least one solution to the PDE on a dense set of points in $S$ and $\partial S$. If $\mathcal{F}u,\mathcal{G}u$ are continuous for functions $u\in\mathcal{H}$, there is a solution to the PDE on all of $S$ and $\partial S$, and our solutions to (6) converge to $\tilde{u}_{\infty}$ where
@@
That is, $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}$ converges to the norm-minimizing solution to the PDE given by (1) and (3). Also note that if at least one strong-form solution to the PDE of (1) and (3) exists and can be extended to a function $u\in\mathcal{H}$, then $u$ is feasible for each (6) and $\left\|\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}\right\|_{\mathcal{H}}\leq\left\|u\right\|_{\mathcal{H}}$ for each $n$, so Proposition 2.1 and Corollary 2.3 apply; $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}\to\tilde{u}_{\infty}$, where $\tilde{u}_{\infty}$ also solves the PDE of (1) and (3).
@@ 2.2 High-Order Bounds on Functions with Scattered Zeros
(See Def. 3.6 from [33]) A domain $U\subset\mathbb{R}^{m}$ satisfies an interior cone condition if there exists an angle $\theta\in\left(0,\frac{\pi}{2}\right)$ and a radius $r>0$ such that for all $\boldsymbol{x}\in U$, there exists a unit vector $\boldsymbol{\xi}\left(\boldsymbol{x}\right)$ such that
@@
(Adapted from Corollary 6 of [31]) Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(U\right)$ where $U\subset\mathbb{R}^{m}$ is bounded and satisfies an interior cone condition, and $m\geq 3$ if $q=0$. Assume $f$ is exactly zero on $\left\{x_{k}\right\}_{k=1}^{\tilde{N}}$ with fill distance $h_{\text{max}}$ on $U$, then for all $p\leq q$, there exist constants $C_{m,p},h_{0,m,p}>0$ such that as long as $h_{\text{max}}<h_{0,m,p}$,
@@
On compact manifolds $S$ with dimension $m$, the same theorem holds, with the sufficient conditions that $S$ can be parametrized by a finite atlas $\left\{\sigma_{k}\right\}_{k=1}^{M}$ with $0<M<\infty$ such that $\sigma_{k}:U_{k}\to S$ is $C^{\left\lceil\frac{m}{2}+q+\frac{1}{2}\right\rceil}$ with bounded derivatives and has a bounded inverse tensor, and each $U_{k}$ is bounded and satisfies an interior cone condition. The theorem may then be applied to $f\circ\sigma_{k}$ on $U_{k}$. The utility of Theorem 2.7 is that if our Hilbert space $\mathcal{H}$ is such that $\left\|\mathcal{F}v\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\leq A\left\|v\right\|_{\mathcal{H}}$ and $\left\|\mathcal{G}v\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(\partial S\right)}\leq B\left\|v\right\|_{\mathcal{H}}$ for each $v\in\mathcal{H}$ and some constants $A,B$, then we can immediately state estimates for $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}$ and $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left(\partial S\right)}$ where $\tilde{u}$ is the solution to (5). In particular, if a solution $u$ to the PDE given by (1) and (3) exists and $\left\|\mathcal{F}v\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\leq A\left\|v\right\|_{\mathcal{H}}$ for each $v\in\mathcal{H}$,
@@
Then, $\left\|\tilde{u}-u\right\|_{\mathcal{H}}\to 0$ by Proposition 2.1, so if we invoke Theorem 2.7 (assuming $S$ satisfies sufficient conditions), noting that $\mathcal{F}\tilde{u}-f$ has scattered zeros, we have, for $p\leq q$,
@@
with a similar expression for $\mathcal{G}\tilde{u}-g$ again. Leaving $f,g$, and $S$ fixed (or $f,g$ varying, but bounded), but varying $h_{\text{max}}$, $\left\|\tilde{u}\right\|_{\mathcal{H}}$ either converges to a non-zero value, assuming $f$ or $g$ are non-zero, or diverges to infinity due to Proposition 2.1. With this in mind, we will often assume for simplicity that the method given by (5) is $p^{\text{th}}$-order in the sense that,
@@
for some constants $A_{p},B_{p}>0$. We again note that (12) holds regardless of whether the PDE given by (1) and (3) is even solvable in $\mathcal{H}$; if it is, then $\left\|\tilde{u}\right\|_{\mathcal{H}}$ is bounded and $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)},$ $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left({\partial S}\right)}\to 0$. If it is not solvable, $\left\|\tilde{u}\right\|_{\mathcal{H}}$ will diverge, potentially faster than $h_{\text{max}}^{-p}$, so $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}$ and $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left({\partial S}\right)}$ could diverge.

### v2

@@ 1 Introduction
For meshfree methods, techniques that rely on knowing the exact integral of a class of functions (such as RBFs) on the domain are typically not available, since these exact integrals are often unknown. Meshfree integration on surface point clouds is therefore often limited to Monte Carlo methods, which have been employed for eigenfunction estimation using the weak form of differential operators [11, 40]. Even though the underlying meshfree interpolation technique used is high-order, Monte Carlo integration is not, which limits the overall rate of convergence. Monte Carlo integration also requires knowledge of the distribution from which a point cloud is sampled, which may not be available, and it may not be straightforward to sample points from a surface uniformly. If the points satisfy a minimum distance condition, so that no two points are too close together, a potential integration technique for non-uniform point clouds is to use local projections to the tangent plane and a triangulation to compute Voronoi cell areas (see [17]), though this has a limited order of convergence. Existing high-order meshfree integration methods on surfaces are often restricted to specific surfaces where exact integrals of RBFs are known, such as the sphere [14] or an explicit basis for a Hilbert space on the surface can be written down, as in Stein reproducing kernel methods from the statistics literature (see [1], for example). Reeger, Fornberg, and Watts [27] have also introduced a high-order ($\mathcal{O}\left(h^{7}\right)$, where $h$ is the point spacing), RBF-based method for general closed surfaces that requires first obtaining a triangulation of the point cloud. This is essentially a hybrid approach between a meshfree and meshed method in that while a triangulation of the point cloud is required, a meshfree method is ultimately used to compute integrals on each surface patch projected to a triangle with high accuracy. A version for surfaces with boundary has also been developed [26].
@@
Our contribution consists of two super-algebraically convergent methods based on the divergence theorem and norm-minimizing Hermite-Birkhoff interpolation schemes. Such interpolation schemes include Hermite RBFs [30, 8] and minimum norm methods using Fourier series [5, 36]. We review these interpolation methods in Section 2, including a generalized convergence statement in Proposition 2.1, a discussion on a general approach to high-order convergence in 2.2, and two examples in 2.3. We then introduce our integration methods in Section 3. Our approaches do not require specific quadrature points (the points may be randomly scattered and non-uniformly selected) and only require the knowledge of basic domain information such as normal vectors at scattered points on the surface. Method 1 can be used to directly evaluate the ratio of two integrals on a closed surface or a surface with boundary; this is useful when a constant multiple of the integral is all that is needed. For example, the method could be used when the average value of a function on the surface is desired. Method 2 requires surfaces to have boundary, but we discuss simple strategies for dividing closed surfaces into multiple surfaces with boundary in Subsection 3.5. Unlike the first method, method 2 estimates integrals themselves rather than a ratio, which can be used to estimate surface area with high precision. We show that these schemes converge at the same rate as the underlying interpolation methods, regardless of the point distribution; this leads to arbitrarily high-order schemes. Then, we test the integration techniques for a variety of problems on flat domains and on surfaces in Section 4. We conclude by introducing a more general view of norm-minimizing meshfree methods suitable for handling singularities, and then adapt our methods for singular integrals in Subsection 4.2; such integrals appear in boundary integral solutions to PDEs. Code for running the numerical tests in this paper is available at https://github.com/venndaniel/meshfree_integration.
@@ 2.1 Norm-Minimizing Hermite-Birkhoff Interpolants
(3) $\mathcal{G}u=g\text{ on }\partial S,$
@@
where $\{\boldsymbol{y}_{j}\}_{j=1}^{\tilde{N}_{\partial}}\subset\partial S$ is a set of points on the boundary of $S$. The goal is then to ensure that $\mathcal{F}\tilde{u}\to f$ on $S$ and $\mathcal{G}\tilde{u}\to g$ on $\partial S$ in some sense as $\tilde{N},\tilde{N}_{\partial}\to\infty$. If the PDE given by (1) and (3) is suitably well-posed, we may then be able to show that $\tilde{u}\to u$, with the manner of convergence ($L^{2},L^{\infty},H^{p}$, pointwise, etc.) depending on the PDE and method chosen (see, for example, Proposition 11 of [36]).
Of course, simply imposing $\left(\mathcal{F}\tilde{u}\right)\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)$ on each $\boldsymbol{x}_{j}$ is not enough to ensure that $\mathcal{F}\tilde{u}\to f$ as more collocation points are added. A classic counterexample is the interpolation of Runge’s function ($f\,(x)=1/(1+25x^{2})$) with polynomials of increasing degree on $\left[-1,1\right]$ and evenly spaced points $\left\{\boldsymbol{x}_{j}\right\}$; this corresponds to an identity operator $\mathcal{F}$ and results in an interpolant $\tilde{u}$ that is unbounded as $\tilde{N}\to\infty$. While basic, this counterexample demonstrates one possible manner in which interpolation can fail: the interpolant can become unbounded. This motivates a constrained approach. For some Hilbert space of functions $\mathcal{H}$ defined on a domain $\Omega\supseteq S$, we solve
@@
For this problem to be uniquely solvable, we require $\mathcal{L}$ to be bounded in the $\mathcal{H}\to\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ sense, which means that both $\mathcal{F}$ and $\mathcal{G}$ evaluated at individual points must be bounded as operators from $\mathcal{H}$ to $\mathbb{C}$. We also need the constraint set to be non-empty for each $\boldsymbol{f}$; $\mathcal{L}$ must be surjective. We note that for reasonable choices of $\mathcal{H}$ (examples are discussed in Subsection 2.3), $\mathcal{L}$ is generally surjective for a finite number of collocation points even if the PDE given by (1) and (3) is not solvable globally. In this case, problem (6) has a few important properties. Let $\tilde{u}\in\mathcal{H}$ be the solution to (6), then $\tilde{u}\in\mathcal{R}\,(\mathcal{L}^{*})=(\mathcal{N}\,(\mathcal{L}))^{\perp}$ (noting that $\mathcal{R}\left(\mathcal{L}^{*}\right)$ is finite-dimensional and therefore closed). We can therefore find $\tilde{u}$ by solving
@@
Note that $\tilde{u}_{n}$ is well-defined due to a standard result in functional analysis (see Theorem 3.3-1 of Kreyszig’s text [15], for example), since $V_{n}$ is non-empty, closed, and convex by assumption. Furthermore, $\left\{\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\right\}_{n=1}^{\infty}$ is non-decreasing, since $V_{n_{1}}\supseteq V_{n_{2}}$ whenever $n_{1}<n_{2}$. For any convex combination $\left(1-\alpha\right)\tilde{u}_{n_{1}}+\alpha\tilde{u}_{n_{2}}$ such that $n_{1}<n_{2}$ and $\alpha\in\left(0,1\right)$,
$\left\|\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}$ $\leq\left\|\alpha\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right)+\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}$
$=\left\|\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}+\alpha^{2}\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}+2\alpha\,\text{Re}\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}},\tilde{u}_{n_{1}}\right)_{\mathcal{H}}$
@@
Finally, $\left\{\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\right\}_{n=1}^{\infty}$ is bounded and non-decreasing, and therefore converges, which implies $\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}\to 0$ as $n_{1},n_{2}\to\infty$ due to (11). The sequence $\left\{\tilde{u}_{n}\right\}_{n=1}^{\infty}$ is therefore Cauchy and must converge to some $\tilde{u}_{\infty}\in\mathcal{H}$, since $\mathcal{H}$ is a Hilbert space. Furthermore, for each $m$, $\left\{\tilde{u}_{n}\right\}_{n=m}^{\infty}\subset V_{m}$, which is closed, so $\tilde{u}_{\infty}\in V_{m}$ as well. Therefore, $\tilde{u}_{\infty}\in\bigcap_{n=1}^{\infty}V_{n}$.
@@
For our optimization problem applied to PDEs, assume that $\{\boldsymbol{x}_{j}\}_{j=1}^{\infty}$ is dense in $S$ and that $\{\boldsymbol{y}_{j}\}_{j=1}^{\infty}$ is dense in $\partial S$. Then, Corollary 2.3 implies that if our solution $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}$ to the optimization problem (6) remains bounded in $\mathcal{H}$ as $\tilde{N},\tilde{N}_{\partial}\to\infty$, there is at least one solution to the PDE on a dense set of points in $S$ and $\partial S$. If $\mathcal{F}u,\mathcal{G}u$ are continuous for functions $u\in\mathcal{H}$, there is a solution to the PDE on all of $S$ and $\partial S$, and our solutions to (6) converge to $\tilde{u}_{\infty}$ where
@@
That is, $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}$ converges to the norm-minimizing solution to the PDE given by (1) and (3). Also note that if at least one strong-form solution to the PDE of (1) and (3) exists and can be extended to a function $u\in\mathcal{H}$, then $u$ is feasible for each (6) and $\|\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}\|_{\mathcal{H}}\leq\left\|u\right\|_{\mathcal{H}}$ for each $n$, so Proposition 2.1 and Corollary 2.3 apply; $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}\to\tilde{u}_{\infty}$, where $\tilde{u}_{\infty}$ also solves the PDE of (1) and (3).
@@ 2.2 High-Order Bounds on Functions with Scattered Zeros
(See Def. 3.6 from [38]) A domain $U\subset\mathbb{R}^{m}$ satisfies an interior cone condition if there exists an angle $\theta\in(0,\frac{\pi}{2})$ and a radius $r>0$ such that for all $\boldsymbol{x}\in U$, there exists a unit vector $\boldsymbol{\xi}\left(\boldsymbol{x}\right)$ such that
@@
(Adapted from Corollary 6 of [36]) Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\,(U)$ where $U\subset\mathbb{R}^{m}$ is bounded and satisfies an interior cone condition, and $m\geq 3$ if $q=0$. Assume $f$ is exactly zero on $\{x_{k}\}_{k=1}^{\tilde{N}}$ with fill distance $h_{\text{max}}$ on $U$, then for all $p\leq q$, there exist constants $C_{m,p},h_{0,m,p}>0$ such that as long as $h_{\text{max}}<h_{0,m,p}$,
@@
On compact manifolds $S$ with dimension $m$, the same theorem holds, with the sufficient conditions that $S$ can be parametrized by a finite atlas $\left\{\sigma_{k}\right\}_{k=1}^{M}$ with $0<M<\infty$ such that $\sigma_{k}:U_{k}\to S$ is $C^{\left\lceil\frac{m}{2}+q+\frac{1}{2}\right\rceil}$ with bounded derivatives and has a bounded inverse tensor, and each $U_{k}$ is bounded and satisfies an interior cone condition. The theorem may then be applied to $f\circ\sigma_{k}$ on $U_{k}$. The utility of Theorem 2.7 is that if our Hilbert space $\mathcal{H}$ is such that $\|\mathcal{F}v\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\leq A\left\|v\right\|_{\mathcal{H}}$ and $\|\mathcal{G}v\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(\partial S\right)}\leq B\left\|v\right\|_{\mathcal{H}}$ for each $v\in\mathcal{H}$ and some constants $A,B$, then we can immediately state estimates for $\|\mathcal{F}\tilde{u}-f\|_{L^{\infty}\left(S\right)}$ and $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left(\partial S\right)}$ where $\tilde{u}$ is the solution to (5).
In particular, if any strong-form solution $u$ to the PDE given by (1) and (3) exists and can be extended to $\mathcal{H}$, and $\left\|\mathcal{F}v\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\leq A\left\|v\right\|_{\mathcal{H}}$ for each $v\in\mathcal{H}$, then
@@
Now, if such a strong-form solution $u$ exists, then solutions to (5) are bounded, $\tilde{u}_{\infty}$ exists by Proposition 2.1, and $\tilde{u}\to\tilde{u}_{\infty}$. If the sets of points $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\infty}$ and $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\infty}$ are dense on $S$ and $\partial S$ and functions in $\mathcal{H}$ are sufficiently smooth, then $\tilde{u}_{\infty}$ restricted to $S$ is also a strong form solution and
$\left\|\mathcal{F}\tilde{u}-f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}$ $\leq A\left\|\tilde{u}-\tilde{u}_{\infty}\right\|_{\mathcal{H}}\to 0.$
Therefore, if we invoke Theorem 2.7 (assuming $S$ satisfies sufficient conditions), noting that $\mathcal{F}\tilde{u}-f$ has scattered zeros, we have, for $p\leq q$,
@@
with a similar expression for $\mathcal{G}\tilde{u}-g$ again. Leaving $f,g$, and $S$ fixed (or $f,g$ varying, but bounded), but varying $h_{\text{max}}$, $\|\tilde{u}\|_{\mathcal{H}}$ either converges to a non-zero value, assuming $f$ or $g$ are non-zero, or diverges to infinity due to Proposition 2.1. With this in mind, we will often assume for simplicity that the method given by (5) is $p^{\text{th}}$-order in the sense that,
@@
for some constants $A_{p},B_{p}>0$. Note that this simplification just relies on the fact that $\left\|\tilde{u}\right\|_{\mathcal{H}}$ is certainly bounded below by a positive constant as long as the right-hand sides of the constraints from (5) are not identically zero, so some constant multiple of $\left\|\tilde{u}\right\|_{\mathcal{H}}$ is greater than $\|f\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}$. We reiterate that $f$ here is either fixed or only allowed to vary boundedly.
We again note that (12) holds regardless of whether the PDE given by (1) and (3) is even solvable in $\mathcal{H}$; if it is, then $\left\|\tilde{u}\right\|_{\mathcal{H}}$ is bounded and $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)},$ $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left({\partial S}\right)}\to 0$. If it is not solvable, $\left\|\tilde{u}\right\|_{\mathcal{H}}$ will diverge, potentially faster than $h_{\text{max}}^{-p}$, so $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}$ and $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left({\partial S}\right)}$ could diverge.